OpenStudy (anonymous):

PLEASE HELP HELP what type of transformation relates y=x squared and y+(x-5) squared

OpenStudy (jdoe0001):

you mean from \(\bf y=x^2\quad to\quad y=(x+5)^2 ?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

instead of to it says and

OpenStudy (confluxepic):

Since it is in parenthesis the line is shifted to the right by 5 units.

OpenStudy (anonymous):

choices are a.horizontal translation b.Vertical translation c. vertical reflection D. reflection in y axis

OpenStudy (anonymous):

so it is a vertical reflection?

OpenStudy (confluxepic):

Horizontal translation.